OpenStudy (anonymous):

Find the product of (3x − 7y)2. 9x2 − 42xy + 49y2 9x2 + 42xy + 49y2 9x2 − 49y2 9x2 + 49y2

HanAkoSolo (jamierox4ev3r):

is it 2(3x-7y) or (3x-7y)^2<----exponenet

OpenStudy (anonymous):

yes @Jamierox4ev3r

OpenStudy (ankit042):

use the formula (a-b)^2=a^2 +b^2-2ab

HanAkoSolo (jamierox4ev3r):

wait but which one lol

HanAkoSolo (jamierox4ev3r):

the first one or the second one?

OpenStudy (anonymous):

second one @Jamierox4ev3r

HanAkoSolo (jamierox4ev3r):

oh okay just picture it like this: (3x-7y)(3x-7y)

HanAkoSolo (jamierox4ev3r):

if you use the property of distribution, you could solve this

HanAkoSolo (jamierox4ev3r):

(3x-7y)(3x-7y) 9x^2-21xy-21xy+49y^2

HanAkoSolo (jamierox4ev3r):

then you combine like terms

HanAkoSolo (jamierox4ev3r):

9x^2+49y^2-42xy

OpenStudy (uri):

(a-b)^2=a^2 +b^2-2ab (3x-7y)^2=(3x)^2+(7y)^2-2(3)(7) =9x^2+49x^2-42xy

HanAkoSolo (jamierox4ev3r):

that is your final answer does that make sense @wendyrodriguez0114 ?

OpenStudy (anonymous):

some what because you combine right? @Jamierox4ev3r

OpenStudy (anonymous):

thank you @Jamierox4ev3r

HanAkoSolo (jamierox4ev3r):

no problem pleasure working with you :DD

HanAkoSolo (jamierox4ev3r):

but if you need it to be written by standard form, you would write your answer the way @Uri wrote it.